# Task Me Anything

### v1

@@ 2.3 Addressing user queries
Fitting similarly samples a random subset and evaluates MLMs. The results are used to train an efficient function approximator for each MLM. This function approximator learns to predict an MLM’s performance on a task, by featurizing the task-metadata—never actually generating the task instance itself. While many model choices are applicable, we adopt the Gaussian Process regressor throughout this work since it renders stable performance in preliminary studies. It uses this function to approximate the MLM’s performance on the remaining task space.
@@
Active is similar to fitting but iteratively trains each function approximator using active learning. Given a smaller subset, it trains an initial function, which is then used to sample the most uncertain task instances. MLMs are evaluated on these uncertain instances; the results are used to re-train the functions again.
@@ 3 Evaluating MLMs using Task-Me-Anything
We adopt the accuracy of the model on a task to capture the model’s performance. However, one task can contain numerous concrete task instances. In practice, we randomly generate $n$ task instances for a task and then use the model’s accuracy on the $n$ task instances as a proxy of the model’s accuracy on the task. For prompts used to evaluate the models, to fairly evaluate the model’s performance and enhance the robustness of the results. We use two versions of prompts: a succinct prompt and a detailed prompt. The succinct version simply adds ’Select from the following choices’ between the question and the options [24], while the detailed prompt includes more instructions such as: ’Based on the image/video", and also enclose the options within parentheses (e.g., “(A) camera (B) telephone”)’ and ends the prompt with ’Best Option: (’ to guide the model to output the option only [53]. The exact prompt template can be found in Figure 4." For option extraction, we match the model output to three types of option representations: 1) option identifier, e.g., “(A)”, 2) option name, e.g., “camera”, and 3) option identifier and name, e.g., “(A) camera” in order to increase the recall of the option extraction.
@@
Task-Me-Anything allows users to query for tasks that most resemble their application. As such, Task-Me-Anything doesn’t have to be limited to a static leaderboard commonly seen with most other benchmarks. Instead, we make Task-Me-Anything’s findings accessible through an interactive graphical user interface. Our interface allows users to specify their needs without writing any code. They can subselect parts of the taxonomy that best represent their application. We use the evaluation results in Task-Me-Anything-DB obtained our explorations to build a simple example interface: Task-Me-Anything-UI  [^1] https://huggingface.co/spaces/zixianma/TaskMeAnything-UI. It consists of four tabs: the overall tab reports model performance across a dozen MLM across different subsets of Task-Me-Anything’s taxonomy; the task embedding tab visualizes different task instances in a 2D space and allows users to observe model behavior across similar tasks; the surprisingness tab highlights tasks where a model achieves surprisingly better or worse performance compared to similar tasks; and the query interface supports users to conduct query-centric investigation of models’ capabilities or limitations using the four types of fine-grained user queries mentioned above (Figure 5). More details can be found in the Appendix D.3.
@@ 4 Validating and ablating Task-Me-Anything
We evaluate the proposed query results approximation algorithms on 1,137 queries across the 4 query types (Table 1). To measure the quality of the approximation, we use the evaluation results from Task-Me-Anything-DB as ground truth query results. From Table 1, we can see that the Active method outperforms both the Random and Fitting methods across nearly all query types, yet there is still room for future improvement. More details of experiments and results are in Appendix F.
@@ 5 Analysing MLMs with Task-Me-Anything
We use Task-Me-Anything to conduct multiple analyses to highlight its different use cases, while simultaneously drawing insights about today’s MLMs (More details in Appendix G). Specifically, we evaluated 18 MLMs on Task-Me-Anything-Random for Query 1 and 4 and reused the evaluation results of Task-Me-Anything-DB for Query 2, 3, 5, and 6. Finally, we leverage Task-Me-Anything to provide an in-depth analysis on GPT4o as Query 7.
@@ 5.1 Query 1: How do models perform over a random subset of all possible questions?
We evaluated 18 MLMs on the Task-Me-Anything-Random test set (Figure 6) to gain an overview of model performance. The detailed prompt typically yields better results; however, certain models, like GPT4V, perform much better with the succinct prompt, indicating that current models are still prompt-sensitive.
For ImageQA tasks, the latest open-sourced models, such as InternVL-Chat-1.5-24B and LLaVA-Next-34B, perform better than popular proprietary models, achieving state-of-the-art performance, which is also shown in recent benchmarking results [16]. Notably, models like InstructBLIP-7B and Qwen-VL perform significantly better with detailed prompt than succinct prompt. For VideoQA tasks, we also evaluated larger or proprietary ImageQA models, like GPT4V, by concatenating four frames of a video into a single picture. Notably, Video-LLaVA-7B perform much better with succinct prompts than other small open-source models.
@@
5.2 Query 2: What skills are MLMs best and worst at?
@@
Further, Task-Me-Anything also enables us to conduct analyses of models’ fine-grained skills such as recognizing a specific type of object, attribute, or relation. For example, on ImageQA tasks, we find that on average models are better at recognizing plants, understanding mood and comprehending spatial relations between real-world objects (Figure 9). Nevertheless, some models might showcase different strengths: LLaVA-13B is better at recognizing animals (Figure 9 (a)), and InstructBLIP-7B is better at understanding emotional relationships (Figure 9 (c)). On the other hand, for VideoQA tasks, we learn that models are better at recognizing vehicles, material and understanding spatial relationships (Figures 10 and 11).
5.3 Query 3: what is the best MLM for each specific skill?
@@
5.4 Query 4: How does the best open-source model compare against the best proprietary model across skills?
@@
5.5 Query 5: How do small models compare against large models?
@@
5.6 Query 6: Are models’ strengths and weaknesses consistent across visual inputs?
@@
5.7 Query 7: What is today’s popular proprietary model, GPT4o, bad at?
@@ C.3 Efficient Query Approximation Algorithms
Figure 21: Summary of task generator card sections and suggested prompts for each. Task generator cards for all the included task generators can be found in Appendix I.
@@ E.2 A breakdown of Table 8
Table 10: 2D sticker image
@@
Table 11: 3D tabletop scene part 1
@@
Table 12: 3D tabletop scene part 2
@@
Table 13: Real images with Scene Graph
@@ E.3 A breakdown of Table 9
Table 14: 3D tabletop scene
@@
Table 15: Real videos with Scene Graph
@@
Appendix F Details of Experiments on Query Results Approximation Algorithms

### v2

@@ 2.3 Addressing user queries
Fitting similarly samples a random subset and evaluates MLMs. The results are used to train an efficient function approximator for each MLM. This function approximator learns to predict an MLM’s performance on a task, by featurizing the task plan—never actually generating the task instance itself. While many model choices are applicable, we adopt the Gaussian Process regressor throughout this work since it renders stable performance in preliminary studies. It uses this function to approximate the MLM’s performance on the remaining task space.
@@
Active is similar to fitting but iteratively trains each function approximator using active learning. Given a smaller subset, it trains an initial function, which is then used to sample the most task instances that are most likely to lead to improved approximation results. MLMs are evaluated on these samples; the results are used to re-train the functions again.
@@ 3 Evaluating MLMs using Task-Me-Anything
We adopt the accuracy of the model on a task to capture the model’s performance. However, one task can contain numerous concrete task instances. In practice, we randomly generate $n$ task instances for a task and then use the model’s accuracy on the $n$ task instances as a proxy of the model’s accuracy on the task. For prompts used to evaluate the models, to fairly evaluate the model’s performance and enhance the robustness of the results, we use two versions of prompts: a succinct prompt and a detailed prompt. The succinct version simply adds ’Select from the following choices’ between the question and the options [24], while the detailed prompt includes more instructions such as: ’Based on the image/video", and also enclose the options within parentheses (e.g., “(A) camera (B) telephone”)’ and ends the prompt with ’Best Option: (’ to guide the model to output the option only [53]. The exact prompt template can be found in Figure 4." For option extraction, we match the model output to three types of option representations: 1) option identifier, e.g., “(A)”, 2) option name, e.g., “camera”, and 3) option identifier and name, e.g., “(A) camera” in order to increase the recall of the option extraction.
@@
Task-Me-Anything allows users to query for tasks that most resemble their application. As such, Task-Me-Anything doesn’t have to be limited to a static leaderboard commonly seen with most other benchmarks. Instead, we make Task-Me-Anything’s findings accessible through an interactive graphical user interface. Our interface allows users to specify their needs without writing any code. They can select parts of the taxonomy that best represent their application. We use the evaluation results in Task-Me-Anything-DB obtained our explorations to build a simple example interface: Task-Me-Anything-UI  [^1] https://huggingface.co/spaces/zixianma/TaskMeAnything-UI. It consists of four tabs: the overall tab reports model performance across a dozen MLM across different subsets of Task-Me-Anything’s taxonomy; the task embedding tab visualizes different task instances in a 2D space and allows users to observe model behavior across similar tasks; the surprisingness tab highlights tasks where a model achieves surprisingly better or worse performance compared to similar tasks; and the query interface supports users to conduct query-centric investigation of models’ capabilities or limitations using the four types of fine-grained user queries mentioned above (Figure 5). More details can be found in the Appendix D.3.
@@ 4 Validating and ablating Task-Me-Anything
We evaluate the proposed query results approximation algorithms on 1,137 queries across the 4 query types (Table 1). To measure the quality of the approximation, we use the evaluation results from Task-Me-Anything-DB as ground truth query results. From Table 1, we can see that the Active method outperforms both the Random and Fitting methods across nearly all query types, yet there is still room for future improvement. More details of experiments and results are in Appendix G.
@@
Comparison with existing benchmarks.
According to Figure 7, we found that there is a strong positive correlation between models’ rankings on Task-Me-Anything and on other benchmarks, with strong correlations ($\geq$0.8) on the most commonly used ones such as MMMU and MMBench. Notably, the average correlation between Task-Me-Anything and the other benchmarks is 0.77, which is greater than that of some other benchmarks such as ScienceQA and LLaVABench (0.70 and 0.57). These results suggest that the evaluation results on our benchmark align with those on existing benchmarks.
Figure 6: Correlations of Task-Me-Anything with other popular MLM benchmarks.
Figure 7: Correlations of synthetic data with real data in Task-Me-Anything.
Analysis on synthetic data in Task-Me-Anything.
Task-Me-Anything leverages both synthetic and real images/videos for evaluating MLM models. While synthetic data are controllable, low-cost, its effectiveness compared to real data remains unclear, particularly given that most benchmarks rely heavily on real datasets. To quantitatively justify the transferability between evaluations on synthetic and real-world images, we compared 18 models’ rankings on synthetic vs. realistic images and obsered strong positive correlations between them. Across the three different image sources in Task-Me-Anything, we found that the correlation between models’ rankings on 2D and 3D synthetic images is the strongest (r=0.99). While the correlations between models’ rankings on real and 2D images, and between real and 3D images are slightly smaller (r=0.85 and 0.79 respectively), these numbers still suggest strong positive correlations between models’ performance on synthetic images and on real images (Figure 7).
@@ 5 Analysing MLMs with Task-Me-Anything
We use Task-Me-Anything to conduct multiple analyses to highlight its different use cases, while simultaneously drawing insights about today’s MLMs (More details in Appendix H). Specifically, we evaluated 18 MLMs on Task-Me-Anything-Random for Query 1 and 4 and reused the evaluation results of Task-Me-Anything-DB for Query 2, 3, 5, and 6. Finally, we leverage Task-Me-Anything to provide an in-depth analysis on GPT4o as Query 7.
@@ 5.1 Query 1: How do models perform over a random subset of all possible questions?
We evaluated 18 MLMs on the Task-Me-Anything-Random test set (Figure 8) to gain an overview of model performance. (Details in Appendix E).
For ImageQA tasks, the latest open-sourced models, such as InternVL-Chat-1.5-24B and LLaVA-Next-34B, surprisingly outperform popular proprietary models, achieving state-of-the-art performance as shown in recent benchmarking results [16].
For VideoQA tasks, in addition to natural VideoQA models, we evaluated larger or proprietary ImageQA models, such as GPT4V, by combining four frames of a video into a single picture. We observed that natural VideoQA models, such as Video-LLaVA-7B, still have a significant performance gap compared to large ImageQA models, even though ImageQA models are not trained on video and only use a combined image rather than the full video.
@@
5.2 Query 2: Are current models sensitive to prompts?
Our evaluation indicates that current models are still prompt-sensitive (Figure 9). We tested our model on two versions of prompts: a succinct version that only describes the question and a detailed version that includes more instructions.
First, we observed that MLMs are still highly prompt-sensitive, and using different prompts might result in over 40% performance differences.
While detailed prompts generally lead to better results, certain models, such as GPT4V, perform notably better with more succinct prompts. In contrast, models like InstructBLIP-7B and Qwen-VL exhibit a marked improvement with detailed prompts compared to succinct ones. (Details in Appendix E).
Figure 9: The absolute difference of model performance in different prompts in Task-Me-Anything-Random.
5.3 Query 3: What skills are MLMs best and worst at?
@@
Further, Task-Me-Anything also enables us to conduct analyses of models’ fine-grained skills such as recognizing a specific type of object, attribute, or relation. For example, on ImageQA tasks, we find that on average models are better at recognizing plants, understanding mood and comprehending spatial relations between real-world objects (Figure 12). Nevertheless, some models might showcase different strengths: LLaVA-13B is better at recognizing animals (Figure 12 (a)), and InstructBLIP-7B is better at understanding emotional relationships (Figure 12 (c)). On the other hand, for VideoQA tasks, we learn that models are better at recognizing vehicles, material and understanding spatial relationships (Figures 13 and 14).
5.4 Query 4: what is the best MLM for each specific skill?
@@
5.5 Query 5: How does the best open-source model compare against the best proprietary model across skills?
@@
5.6 Query 6: How do small models compare against large models?
@@
5.7 Query 7: Are models’ strengths and weaknesses consistent across visual inputs?
@@
5.8 Query 8: What is today’s popular proprietary model, GPT4o, bad at?
@@
5.9 Which models should I use?
When considering which models to use for your tasks, proprietary models generally offer superior performance due to their extensive training on diverse datasets and robust architectures. For tasks requiring high accuracy in object recognition, relation understanding, or attribute identification, GPT4o is currently among the best options. Its ability to handle complex queries and produce refined outputs makes it a top choice for applications demanding reliability and depth.
For those seeking open-source alternatives, LLaVA-Next-34B and InternVL-Chat-1.5-24B stand out as strong contenders, often achieving performance levels comparable to proprietary models while being much smaller in size (20-30B) and capable of running on a single A100 GPU or two A6000 GPUs. In VideoQA tasks, current models are still in their early stages and may underperform compared to popular ImageQA models that only takes several video frames instead of full videos. However, for research and study purposes, Video-LLaVA-7B are currently the state-of-the-art for these tasks.
6 Task-Me-Anything-2024 Benchmark
Finally, we introduce Task-Me-Anything-2024, a benchmark specifically designed to highlight tasks that popular MLMs are still struggling with. With Task-Me-Anything system, we are able to automatically identify the tasks challenging for given MLMs within a given budget. The resultant Task-Me-Anything-2024 provides a comprehensive picture of the limitations of current MLMs.
6.1 Task-Me-Anything-2024 Generation Process.
We conducted Top-K queries with active approximation algorithms on popular open-source MLMs across all task types. For each task type, we use a budget of 300 model inferences for each model to query the worst-performing tasks.
In particular, for ImageQA, we use the following models: InstructBLIP-7B, InstructBLIP-13B, Qwen-VL, Qwen-VL-Chat, LLaVA-7B, LLaVA-13B, GLM-4v, CogVLM2-19B, Idefics2-8B, Phi-3-vision-3B, PaliGemma-3B, and InternVL-Chat-1.5-24B; for VideoQA, we use: Video-ChatGPT-7B, Video-LLaVA-7B, VideoChat2-7B, Video-LLaMA-2-7B, Video-LLaMA-2-13B, Chat-UniVi-7B, Chat-UniVi-13B, and InternVL-Chat-1.5-24B (by taking combined frames).
For each task type (e.g., 3d-how-many), we collected the Top 10 worst-performing tasks for each model in succinct and detailed prompts. By combining the worst-performing tasks of each model, we obtained 12,270 ImageQA and 3,567 VideoQA questions that current popular MLMs are struggling with as our Task-Me-Anything-2024 benchmark.
6.2 Results and analysis.
Figure 22: Model’s overall performance on Task-Me-Anything-2024 and averaged model’s performance drop from Task-Me-Anything-Random to Task-Me-Anything-2024.
We then evaluate open-sourced models and GPT4o on Task-Me-Anything-2024 dataset (Figure 22). Comparing the models’ performance on Task-Me-Anything-Random and Task-Me-Anything-2024, we observed a 10-30% performance drop for each model. Importantly, GPT4o, which wasn’t involved in the Task-Me-Anything-2024 generation process, also performed significantly worse. This validates that Task-Me-Anything-2024 is more challenging than Task-Me-Anything-Random, and our generation process indeed found the questions where popular visual models broadly suffer. Additionally, we observe that the relative performance of models still holds on Task-Me-Anything-2024. Larger or proprietary models consistently perform better than smaller ones.
@@ 8 Conclusion
Acknowledgement
This project was partially funded by Toyota Motor Corporation and OpenAI Superalignment Fellowship.
@@ C.3 Efficient Query Approximation Algorithms
Figure 25: Summary of task generator card sections and suggested prompts for each. Task generator cards for all the included task generators can be found in Appendix J.
@@ E.2 A breakdown of Table 8
Table 10: random-2D sticker image
@@
Table 11: random-3D tabletop scene part 1
@@
Table 12: random-3D tabletop scene part 2
@@
Table 13: random-Real images with Scene Graph
@@ E.3 A breakdown of Table 9
Table 14: random-3D tabletop scene
@@
Table 15: random-Real videos with Scene Graph
@@
Appendix F Details of Model Performance on TaskMeAnything 2024 benchmark
In this section, we present the full results of our evaluation on Task-Me-Anything-2024 with 18 MLMs.
F.1 Raw results of Figure 22
Table 16: 2024-ImageQA. The model performance on 2024 subsets of ImageQA tasks using both the detailed prompt and the succinct prompt. Numbers in parentheses are the number of task instances for each set.
2D sticker image 3D tabletop scene Scene Graph
(3,279) (7,095) (1,896)
InstructBLIP-7B 22.92 1.12 24.94 0.42 43.23 0.38
InstructBLIP-13B 22.20 0.72 21.55 0.45 45.35 0.86
Qwen-VL 23.48 12.02 23.77 12.36 42.68 11.26
Qwen-VL-Chat 25.53 24.55 24.21 25.72 53.04 52.38
LLaVA-7B 25.83 26.09 23.64 23.01 42.67 46.75
LLaVA-13B 27.03 24.85 26.25 25.78 51.01 49.10
GLM-4v 32.76 33.80 37.09 37.33 47.80 44.22
CogVLM2-19B 29.99 31.11 37.29 39.66 49.46 48.25
Idefics2-8B 30.50 32.20 35.32 34.80 44.40 46.61
Phi-3-vision-3B 30.44 32.95 27.88 29.52 46.57 49.46
PaliGemma-3B 32.14 33.77 28.95 29.42 58.45 58.71
InternVL-Chat-1.5-24B 36.78 37.08 43.53 42.55 57.08 52.01
GPT4o 31.46 40.87 31.86 47.46 52.03 53.87
Table 17: 2024-VideoQA. The model performance on 2024 subsets of VideoQA tasks using both the detailed prompt and the succinct prompt. Numbers in parentheses are the number of task instances for each set.
3D tabletop scene Scene Graph
(2,394) (1,173)
Detailed prompt Succinct prompt Detailed prompt Succinct prompt
Video-ChatGPT-7B 12.86 10.87 15.56 13.87
Video-LLaVA-7B 19.59 22.14 21.68 34.54
VideoChat2-7B 21.69 17.42 30.31 19.01
Video-LLaMA-2-7B 18.79 10.59 27.32 20.36
Video-LLaMA-2-13B 17.31 12.34 23.17 15.10
Chat-UniVi-7B 16.48 14.47 36.44 26.66
Chat-UniVi-13B 17.84 15.36 27.30 20.60
InternVL-Chat-1.5-24B 23.67 23.58 54.04 38.02
GPT4o 26.96 34.53 57.88 58.23
F.2 A breakdown of Table 16
Table 18: 2024-2D sticker image
how many what what attribute where where attribute
DP SP DP SP DP SP DP SP DP SP
InstructBLIP-7B 25.25 0.00 16.74 0.00 28.77 0.00 22.75 2.15 21.07 3.46
InstructBLIP-13B 22.11 0.00 14.83 0.00 33.33 0.00 20.60 1.57 20.13 2.04
Qwen-VL 23.27 4.13 21.15 7.64 35.16 11.42 19.74 18.03 18.08 18.87
Qwen-VL-Chat 28.88 21.12 23.49 24.52 36.53 35.31 21.32 24.03 17.45 17.77
LLaVA-7B 29.04 31.19 19.53 23.20 31.20 29.07 26.90 25.75 22.48 21.23
LLaVA-13B 30.86 34.32 21.73 21.44 40.03 34.40 24.46 19.46 18.08 14.62
GLM-4v 43.40 45.05 29.37 31.28 42.31 44.44 27.18 27.18 21.54 21.07
CogVLM2-19B 41.09 43.23 26.87 27.61 42.16 44.75 23.03 23.61 16.82 16.35
Idefics2-8B 42.90 41.91 24.96 26.28 41.70 39.88 25.04 29.33 17.92 23.58
Phi-3-vision-3B 38.94 40.26 27.61 26.28 40.33 44.14 23.32 27.18 22.01 26.89
PaliGemma-3B 33.66 33.83 22.61 24.96 38.81 45.36 37.34 37.48 28.30 27.20
InternVL-Chat-1.5-24B 38.78 47.85 34.95 30.40 48.86 40.33 33.48 34.91 27.83 31.92
GPT4o 36.80 48.35 31.57 41.56 37.29 48.25 28.04 36.77 23.58 29.40
Table 19: 2024-3D tabletop scene part 1
how many what what attribute where where attribute
DP SP DP SP DP SP DP SP DP SP
InstructBLIP-7B 26.30 0.00 17.36 0.00 28.79 0.00 26.31 1.05 25.45 1.81
InstructBLIP-13B 20.49 0.00 15.35 0.16 29.89 0.00 21.23 0.75 21.35 1.15
Qwen-VL 30.28 6.88 20.00 5.74 33.49 14.55 20.33 14.05 20.85 11.82
Qwen-VL-Chat 34.86 32.72 19.84 20.47 36.46 35.84 18.24 18.68 16.09 20.53
LLaVA-7B 21.87 21.87 21.24 24.65 29.58 28.17 24.96 20.48 24.96 19.87
LLaVA-13B 29.66 34.25 23.88 24.50 34.90 32.55 22.87 19.43 21.02 19.54
GLM-4v 48.93 48.78 39.69 38.45 42.88 43.51 24.51 22.12 19.54 19.38
CogVLM2-19B 37.77 44.50 29.77 30.23 44.29 46.17 21.23 24.07 21.02 20.36
Idefics2-8B 42.05 41.13 25.89 24.50 40.85 38.97 29.75 28.85 27.09 29.06
Phi-3-vision-3B 40.98 42.05 25.89 24.65 37.25 38.50 20.03 20.63 18.88 23.65
PaliGemma-3B 37.46 30.58 22.95 26.36 38.03 42.25 33.78 32.14 26.27 26.44
InternVL-Chat-1.5-24B 43.12 49.08 39.84 38.14 50.23 45.54 34.08 33.03 33.17 26.93
GPT4o 32.26 54.43 27.29 40.93 36.62 52.27 22.12 32.59 19.38 26.93
Table 20: 2024-3D tabletop scene part 2
DP SP DP SP DP SP DP SP DP SP DP SP
InstructBLIP-7B 17.26 0.00 31.94 0.00 31.71 0.00 13.04 0.00 27.99 1.78 28.17 0.00
InstructBLIP-13B 16.92 0.17 24.50 0.62 29.20 0.00 9.33 0.15 20.06 1.94 28.76 0.00
Qwen-VL 17.78 8.03 23.88 16.59 28.17 11.50 16.59 11.41 21.68 17.80 28.47 17.55
Qwen-VL-Chat 19.66 18.63 17.36 21.09 31.86 31.12 20.30 22.22 19.90 26.21 31.71 35.40
LLaVA-7B 22.91 23.42 22.17 22.17 26.11 25.22 17.04 19.85 22.98 24.60 26.25 22.86
LLaVA-13B 23.93 25.64 19.38 17.05 32.60 30.24 23.85 24.59 21.84 25.24 34.81 30.53
GLM-4v 35.90 36.75 25.89 25.27 34.96 40.27 55.26 55.26 32.36 31.39 48.08 49.41
CogVLM2-19B 42.39 44.79 26.51 32.25 50.29 52.51 54.52 54.07 29.77 32.20 52.65 55.16
Idefics2-8B 31.79 30.26 22.79 24.50 42.18 43.81 34.96 33.48 37.22 36.57 53.98 51.62
Phi-3-vision-3B 30.43 28.72 15.66 19.38 31.27 34.22 28.59 28.74 27.35 30.10 30.38 34.07
PaliGemma-3B 17.44 21.37 25.27 18.29 30.38 31.56 16.00 19.70 31.88 33.33 38.94 41.59
InternVL-Chat-1.5-24B 44.96 26.84 33.95 41.86 46.76 48.53 50.37 51.85 43.85 43.53 58.55 62.68
GPT4o 38.97 57.61 19.84 38.91 42.04 54.72 32.59 54.07 28.80 43.04 50.59 66.52
Table 21: 2024-Real images with Scene Graph
what attribute what object what relation
DP SP DP SP DP SP
InstructBLIP-7B 42.48 0.00 39.97 0.32 47.25 0.81
InstructBLIP-13B 46.98 0.47 50.55 0.16 38.51 1.94
Qwen-VL 43.10 1.24 46.76 6.48 38.19 26.05
Qwen-VL-Chat 50.54 45.27 47.08 50.87 61.49 61.00
LLaVA-7B 47.44 36.59 54.03 48.97 26.54 54.69
LLaVA-13B 41.71 34.26 54.19 49.61 57.12 63.43
GLM-4v 42.17 40.78 56.56 54.03 44.66 37.86
CogVLM2-19B 45.12 46.51 65.72 65.40 37.54 32.85
Idefics2-8B 38.29 36.12 48.97 53.55 45.95 50.16
Phi-3-vision-3B 45.58 44.03 54.98 60.98 39.16 43.37
PaliGemma-3B 48.84 49.15 74.57 74.88 51.94 52.10
InternVL-Chat-1.5-24B 47.75 42.64 65.24 55.29 58.25 58.09
GPT4o 39.69 41.86 51.50 52.76 64.89 66.99
F.3 A breakdown of Table 17
Table 22: 2024-3D tabletop scene
DP SP DP SP DP SP DP SP DP SP DP SP
Video-ChatGPT-7B 21.54 13.33 10.12 7.65 5.88 7.11 12.41 12.17 21.39 19.40 5.82 5.56
Video-LLaVA-7B 24.36 32.05 20.49 22.22 17.65 18.38 18.25 19.22 16.92 17.66 19.84 23.28
Video-LLaMA-2-7B 22.31 15.38 24.20 12.84 13.73 11.52 13.14 9.25 17.41 8.46 21.96 6.08
Video-LLaMA-2-13B 26.41 15.90 21.98 15.80 13.97 9.56 13.14 12.17 11.69 8.46 16.67 12.17
Chat-UniVi-7B 16.67 14.62 16.54 14.07 13.73 15.93 17.03 15.33 16.42 14.68 18.52 12.17
Chat-UniVi-13B 24.62 23.08 19.51 15.56 12.75 9.80 14.84 17.52 18.41 7.71 16.93 18.52
InternVL-Chat-1.5-24B 35.38 35.90 36.30 36.79 16.67 12.99 18.00 15.82 13.43 16.17 22.22 23.81
GPT4o 31.54 37.69 39.01 53.33 22.06 17.89 26.52 33.82 15.92 26.62 26.72 37.83
Table 23: 2024-Real videos with Scene Graph
what action what object what relation
DP SP DP SP DP SP
Video-ChatGPT-7B 18.40 19.47 13.44 12.66 14.84 9.49
Video-LLaVA-7B 23.47 37.33 16.28 27.13 25.30 39.17
Video-LLaMA-2-7B 21.87 14.13 28.94 24.81 31.14 22.14
Video-LLaMA-2-13B 25.60 16.27 26.87 17.83 17.03 11.19
Chat-UniVi-7B 40.00 28.53 33.07 18.60 36.25 32.85
Chat-UniVi-13B 30.67 20.00 27.13 19.90 24.09 21.90
InternVL-Chat-1.5-24B 63.73 41.07 47.29 28.94 51.09 44.04
GPT4o 57.33 59.73 41.86 39.79 74.45 75.18
Appendix G Details of Experiments on Query Results Approximation Algorithms